Ali walks 680m south 780m east 80m south find the distance and the displacement?
1
Expert's answer
2020-03-16T13:17:56-0400

"S=S_1+S_2+S_3",

where, "S" - overall traveled distance, "S_1 = 680 \; m" - the first path segment to the south, "S_2 = 780 \; m" - path segment to the east, "S_3 = 80 \; m" - the second path segment to the south.

"S=680+780+80=1540 \; m = 1.54 \; km"

"D=\sqrt{D_s^2 + D_e^2}",

where "D" - overall displacement, "D_s=S_1 +S_3= 680+80=760 \; m" - overall displacement to the south, "D_e=S_2=780 \; m" - displacement to the east.

"D=\sqrt{760^2 + 780^2}=\sqrt{577600 + 608400}=\sqrt{118000} \approx 1089 \; m \approx 1.09 \; km"